# DistPred: A Distribution-Free Probabilistic Inference Method for Regression and Forecasting

### v1

@@ B.1 Commonly used TS datasets
The information of the experiment datasets used in this paper are summarized as follows: (1) Electricity Transformer Temperature (ETT) dataset Zhou et al. (2021a), which contains the data collected from two electricity transformers in two separated counties in China, including the load and the oil temperature recorded every 15 minutes (ETTm) or 1 hour (ETTh) between July 2016 and July 2018. (2) Electricity (ECL) dataset  [^1] https://archive.ics.uci.edu/ml/datasets/ElectricityLoadDiagrams20112014 collects the hourly electricity consumption of 321 clients (each column) from 2012 to 2014. (3) Exchange Lai et al. (2018) records the current exchange of 8 different countries from 1990 to 2016. (4) Traffic dataset  [^2] http://pems.dot.ca.gov records the occupation rate of freeway system across State of California measured by 861 sensors. (5) Weather dataset  [^3] https://www.bgc-jena.mpg.de/wetter records every 10 minutes for 21 meteorological indicators in Germany throughout 2020. (6) Solar-Energy Lai et al. (2018) documents the solar power generation of 137 photovoltaic (PV) facilities in the year 2006, with data collected at 10-minute intervals. (7) The PEMS dataset Liu et al. (2022a) comprises publicly available traffic network data from California, collected within 5-minute intervals and encompassing 358 attributes. (8) Illness (ILI) dataset  [^4] https://gis.cdc.gov/grasp/fluview/fluportaldashboard.html describes the influenza-like illness patients in the United States between 2002 and 2021, which records the ratio of patients seen with illness and the total number of the patients. The detailed statistics information of the datasets is shown in Table 4.
The dataset information in terms of their size and number of features is summarized in Table 5.
@@
Table 6: Batch size settings of UCI regression tasks across different models.
PBP MC Dropout Deep Ensembles GCDS CARD (ours)
Boston $32$ $32$ $32$ $32$ $32$
Concrete $32$ $32$ $32$ $32$ $32$
Energy $32$ $32$ $32$ $32$ $32$
Kin8nm $64$ $32$ $64$ $64$ $64$
Naval $64$ $32$ $64$ $64$ $64$
Power $64$ $64$ $64$ $64$ $64$
Protein $100$ $256$ $100$ $256$ $256$
Wine $32$ $32$ $32$ $32$ $32$
Yacht $32$ $32$ $32$ $32$ $32$
Year $256$ $256$ $100$ $256$ $256$
B.2 Distribution free vs. distribution related
@@
Then, we can train DistPred utilizing Equation 13 as the loss function. Currently, DistPred exclusively provides the mean and variance of the response variable.
@@
Appendix C Implementation details
Appendix D Univariate time series forecasting
@@
Appendix E Limitations of DistPred
DistPred is not without its limitations. Firstly, the model is not designed to handle categorical data, and as such, it is not suitable for classification tasks. Secondly, the model is not designed to handle time series data with missing values. Thirdly, the model is not designed to handle multivariate time series data. Finally, the model is not designed to handle time series data with irregular time intervals.

### v2

@@ A.2 Implementation Details
Appendix B Comparisons of Model Efficiency in Time Series Forecasting
We evaluate the inference time and size of DistPred (mean only) in comparison to other Transformer-based models, as shown in Fig. 10. It can be observed that DistPred’s runtime, model size, and memory usage do not significantly increase compared to other models.
Figure 10: QQ-plots and histograms of response variable.
Appendix C Distribution Free v.s. Distribution Related
@@ Appendix B Comparisons of Model Efficiency in Time Series Forecasting
Then, we can train DistPred utilizing Eq. 16 as the loss function. As shown in Table 8, DistPred exclusively provides the mean and variance of the response variable.
@@
Appendix D DistPred for Time Series with Missing Values
The proposed DistPred is equipped to handle time series with missing values scenarios, as both imputation and forecasting are done in a similar manner. Our experiments have demonstrated that DistPred remains SOTA even when a substantial portion of the data is missing. Specifically, in cases where 80% of the PhysioNet2012 data points were absent, the model still maintained competitive performance, as shown in Table 9.
Table 9: Comparison of DistPred with other SOTA models when time series with missing values.
Model M-RNN GP-VAE BRITS USGAN CSDI TimesNet Transformer SAITS DistPred
MSE 0.864$\pm$0.002 0.433$\pm$0.011 0.325$\pm$0.002 0.306$\pm$0.001 0.260$\pm$0.05 0.272$\pm$0.006 0.225$\pm$0.002 0.218$\pm$0.002 0.204$\pm$0.002
MAE 0.674$\pm$0.001 0.4$\pm$0.007 0.246$\pm$0.001 0.25$\pm$0.001 0.211$\pm$0.003 0.266$\pm$0.007 0.209$\pm$0.002 0.202$\pm$0.002 0.196$\pm$0.002
Appendix E Univariate Time Series Forecasting